A recent study shows that , the annual cost of maintaining a building in Accra averages 200 Cedis with a variance of 260 Cedis. If a tax of 30% is introduced on all items associated with the maintenance of building that is, everything made 30% more expensive . Calculate the standard deviation of the annual cost of making a building in Accra.
1
Expert's answer
2020-05-15T17:05:17-0400

"\text {Let Y represents the annual cost of maintaining}\\
\text { a building in Accra, then, E(Y)=200, V(Y)=260,}\\
\text {Let U represents the annual cost of maintaining}\\
\text { a building in Accra after adding a tax of 30\%,}\\
U=1.3Y\\
V(U)=V(1.30Y)=1.30^{2}V(Y)\\
=1.30^{2}(260)=439.4,\\
\therefore \sigma_{U}=\sqrt{V(X)}=\sqrt{439.4}=20.96"